# A Personalized Predictive Model that Jointly Optimizes Discrimination and Calibration

### v1

@@ 1 Introduction
Models developed in health prediction, including for mortality, are generally good for the average individual, but are not tailored to individuals with unique characteristics. This is the motivation behind personalized predictive models (PPMs), where predictions are generated for an index patient using a model developed on a similar subpopulation. This area of research is especially valuable in the context of intensive care unit (ICU) prediction models, where it has been suggested that due to the complexity and ambiguity of illnesses in the ICU, a personalized approach should be adopted to both research and practice [1].
@@
It has been shown that PPMs can lead to better prediction compared to one-size-fits-all models [2, 3]. Consequently, many researchers have proposed algorithms that define a subpopulation of patients that are similar to an index patient $\mathrm{I}$ and train a model using only this subpopulation, ultimately to make a prediction for patient $I$. The most significant differences in these various algorithms are the similarity criterion that is used and the performance measure applied to evaluate the final models.
@@
One group of algorithms used when fitting PPMs, as noted in a scoping review [4], are neighbourhood-based algorithms, in which a group of patients similar to patient $I$ is identified to train the PPM.
@@
Hielscher et al.’s algorithm for the classification of epidemiological data on the hepatic steatosis disorder is an example of a neighbourhood-based algorithm for fitting a PPM [5]. They use a weighted $k$-nearest neighbour ($k$NN) classifier with Euclidean distance to measure similarity, and choose $k$ to be the value that results in the best average sensitivity, specificity, and accuracy after a two-fold validation. Another neighbourhood-based algorithm, proposed by Ng et al, uses a Locally Supervised Metric Learning (LSML) similarity metric, which is based on the Mahalanobis distance measure [6]. To assess the performance of their resulting PPM, they use the area under the receiver operating characteristic curve (AUROC), a commonly calculated measure of model discrimination. A more recent paper by Wang et al. also uses Mahalanobis distance to define similarity between patients in their neighbourhood-based algorithm [7]. They choose the optimal size of subpopulation based on the AUROC, cross-entropy loss and the F1-score, which is a measure of the model’s accuracy.
@@
Non-distance-based similarity metrics used in neighbourhood-based algorithms include a similarity measure proposed by Campillo et al. which uses the logical OR-operator [8]. This measure incorporates weights that need to be learned before they are applied to the features. Then, the similarity measure is used to define a subpopulation consisting of the top $M$ similar patients to an index patient, where $M$ is chosen to maximize the AUROC. Another example is the cosine-similarity measure used in an algorithm proposed by Lee et al. (2015), where the measure is used to define a subpopulation consisting of the top $M$ similar patients to the index patient, where, again, $M$ is chosen to maximize the AUROC. In later work done by Lee, he suggests that choosing the optimal $M$ value will become important when taking these methods into practice [9].
@@
As mentioned above, a common quantity in choosing the subpopulation size $M$ is the AUROC [4, 10], a measure of model discrimination. Discrimination is the ability of the predictive model to distinguish between outcome classes. Another measure of model performance is calibration, which is the agreement between the observed outcomes and the predictions. For prediction models in general, it has been criticized that calibration is not assessed nearly as often as discrimination [11]. This is an issue because poorly calibrated clinical algorithms can be misleading and potentially harmful when making decisions. Van Calster et al. also argue that even when calibration is assessed, it is usually done via external validation after the model has been developed via internal validation. Motivated by this issue, Jiang et al. propose a prediction algorithm that jointly optimizes the discrimination and calibration of a support vector machine by manipulating the decomposition of an overall performance measure, the Brier Score [12]. The Brier Score is a summary measure that is optimized when the correct probabilities are used, making it a proper scoring rule [13]. For models that have similar discrimination, a well-calibrated model has a better Brier Score than a miscalibrated model. However, in a recent paper by Assel et al., they claim that the Brier Score does not appropriately evaluate the clinical utility of prediction models [14].
@@
The remainder of the paper will be structured as follows. In Section 2, we propose an algorithm inspired by Lee et al. to develop a PPM that jointly optimizes both discrimination and calibration. We do this by finding the optimal size of the subpopulation, $M$, that jointly optimizes these performance measures through the use of a proposed loss function, which is an extension of the Brier Score. We use a two-term decomposition of the Brier Score, along with a mixture term, to allow for the flexibility of emphasizing one performance measure over another when training the PPM to find the optimal $M$. Section 3 contains results from a set of simulation studies which show the relationships between $M$ and various performance measures in addition to a sensitivity analysis on the mixture term of our defined loss function. In Section 4, our proposed algorithm is tested on data from the eICU Collaborative Research Database, a large multi-center critical care database from hospitals across the United States, to predict mortality of patients with diseases of the circulatory system [15].
@@ 2.1 Notation
Let there be $N$ patients in the training data, where each patient $k$ is represented by a Euclidean vector, $X_{k}$, $k$ = 1, …, $N$, in the multi-dimension feature space defined by $p$ features. We let $x_{j}$, $j=1,...,p$, represent the $n\times 1$ vector consisting of each patient’s value of a given feature. Further, let $X_{I}$ be the Euclidean vector representation of the index patient, for whom we would like to make a prediction. Note that the index patient is not included in the training dataset.
The focus of the prediction task in this paper is on binary classification. Thus, we consider a binary outcome. The actual outcome and predicted probability of the outcome for a patient $k$ are represented by $y_{k}\in\{0,1\}$ and $p_{k}\in[0,1]$, respectively. Finally, let $M$ be a scalar representing the size of the subpopulation in the training data consisting of similar patients to the index patient.
@@ 2.2 Patient Similarity Metric
We use the cosine similarity metric as it is bounded, interpretable, and able to handle non-continuous data. Further, it is hypothesized that the cosine metric may be more sensitive to details in patient differences when measuring similarities [16]. The cosine similarity between patients $I$ and $k$ is shown in Equation (1) where $\bullet$ and |||||$\cdot$||| | ⋅ | | represent the dot product and the Euclidean vector magnitude, respectively. We label this metric as $CSM(X_{I},X_{k})$.
$CSM(X_{I},X_{k})=\frac{X_{I}\bullet X_{k}}{||X_{I}||||X_{k}||}$ (1)
Equation (1) is a bounded measure of similarity, with $-1\leq CSM(X_{I},X_{k})\leq 1$. $CSM(X_{I},X_{k})=-1$ corresponds to minimum (opposite) similarity, $CSM(X_{I},X_{k})=1$ corresponds to maximum similarity, and a value of 0 refers to orthogonal similarity.
@@ 2.3 Performance Measures
To measure the performance of a PPM, one should measure both the discrimination and calibration of the model [17]. Often, overall performance measures such as the Brier Score, shown in Equation (2), are used, as the score is affected by both model discrimination and calibration.
$BrS=\frac{1}{N}\sum_{k=1}^{N}(y_{k}-p_{k})^{2}$ (2)
In a binary setting, such as the one we consider in this project, the Brier Score can range from 0 for a perfect model to 0.25 for a non-informative model when the outcome proportion is 50%. However, the range of the Brier Score changes depending on the outcome proportion, thus making the interpretation of the Brier Score a function of that proportion. A solution to this problem is to scale the Brier Score by its maximum possible score for a realized outcome proportion under a non-informative model [18]. However, in this work, we do not consider the scaled Brier Score as the interpretation of the Brier Score is unnecessary for our purposes, as will be demonstrated below.
Even when predictive models have high discrimination, the calibration may still be poor due to statistical overfitting [11]. One limitation of the Brier Score is that it may select a model that is miscalibrated with good discrimination over a model that is better calibrated with average discrimination [14]. Thus, it may be advantageous to evaluate the discrimination and calibration separately to obtain a more comprehensive assessment of the predictive performance of the model.
Van Calster et al. have defined a hierarchy of four increasingly strict levels of calibration, referred to as mean, weak, moderate, and strong calibration [19]. We will be using mean, weak, and moderate measures of calibration, as strong calibration is the utopic goal but unrealistic in practice.
Calibration-in-the-large is a measure of mean calibration and evaluates whether the observed event rate is equal to the average predicted probabilities, shown in Equation (3). A value of 0 indicates perfect mean calibration.
$CITL=\frac{1}{N}\sum_{k=1}^{N}|y_{k}-p_{k}|$ (3)
@@
The Integrated Calibration Index (ICI), valued between 0 and 1, is a measure of moderate calibration [20], and is the weighted difference between observed and predicted probabilities, where the observations are weighted by the empirical density function of the predicted probabilities. This measure is used to compare the relative calibration of different prediction models, where a lower value of the ICI corresponds to better calibration.
@@ 2.4 Mixture Loss Function
$L=\frac{\alpha}{N}\sum_{k=1}^{N}(x_{k}-p_{k})(1-2p_{k})+\frac{1-\alpha}{N}\sum_{k=1}^{N}p_{k}(1-p_{k}).$ (4)
The first and second terms are related to calibration and discrimination, respectively. Thus, for larger values of $\alpha$, greater emphasis is put on calibration, while for smaller values of $\alpha$, greater emphasis is placed on discrimination. When $\alpha=0.5$, the loss function is proportional to the Brier Score.
@@ 2.5 Personalized Predictive Modelling Algorithm
We propose an algorithm which sets up a validation procedure to tune the subpopulation size, $M$, for a given value of $\alpha$. We include an external validation step in the algorithm where we estimate the performance of the resulting PPM along with quantifying the uncertainty of these estimates. Step-by-step details of the algorithm are given below.
@@
      Split the TrTe patients into training and testing sets using $K$-fold cross-validation (CV).
@@
        For index patient $I$ in the test set, calculate $d(X_{I},X_{k})$, for $k=1,...,n_{k}$ in the training set, where $d(X_{I},X_{k})$ is the similarity metric between the index patient and the $k^{th}$ training set patient.
@@
          Sort all $d(X_{I},X_{k})$ in descending order, and the patients corresponding to the top $M$ scores form the subpopulation.
@@
          Apply a pre-determined weighting function (eg. uniform, (half-)tricube) to each patient in the subpopulation. Note that the (half-)tricube weights give those more similar to the index patients a greater weight.
@@
      Draw a bootstrap sample from the validation sample (same size as the validation sample).
@@
        For index patient $I$ in the test set, calculate $d(X_{I},X_{k})$, for $k=1,...,n_{k}$ in the training set, where $d(X_{I},X_{k})$ is the similarity metric between the index patient and the $k^{th}$ patient.
@@
    Repeat the entire process $Z$ times, ie. repeat the process with $Z$ different validation holdout samples (and, thus, TrTe samples) to view the consistency of the tuning of $M$ (step (1) above) as well as the model performance measures in different hold-out validation samples. We suggest $Z$ to be approximately ten, though five may be sufficient for some problems.
The proposed algorithm above is done for a given $\alpha$, perhaps one that is determined ahead of time, such as $\alpha=0.5$, where both discrimination and calibration are equally balanced in Equation (4). However, there may be some situations where we want to investigate the performance and resulting $M$ for different scenarios, say a bit more emphasis being placed on calibration over discrimination. We could then implement the above algorithm for a few different $\alpha$ values, perhaps to see how the performance will do as a function of both $\alpha$ and $M$. We will see some of this work in Section 3 and 4 below.

### v2

@@ 1 Introduction
Models developed in health prediction, including for mortality, are generally good for the average individual, but are not tailored to individuals with unique characteristics. This is the motivation behind personalized predictive models (PPMs), where predictions are generated for an index patient using a model developed on a similar subpopulation. This area of research is especially valuable in the context of intensive care unit (ICU) prediction models, where it has been suggested that due to the complexity and ambiguity of illnesses in the ICU, a personalized approach should be adopted to both research and practice [Maslove et al., 2017].
@@
It has been shown that PPMs can lead to better prediction compared to one-size-fits-all models [Celi et al., 2012, Lee et al., 2015]. Consequently, many researchers have proposed algorithms that define a subpopulation of patients that are similar to an index patient $\mathrm{I}$ and train a model using only this subpopulation, ultimately to make a prediction for patient $I$. The most significant differences in these various algorithms are the similarity criterion that is used and the performance measure applied to evaluate the final models.
@@
One group of algorithms used when fitting PPMs, as noted in a scoping review [Sharafoddini et al., 2017], are neighbourhood-based algorithms, in which a group of patients similar to patient $I$ is identified to train the PPM.
@@
Hielscher et al.’s algorithm for the classification of epidemiological data on the hepatic steatosis disorder is an example of a neighbourhood-based algorithm for fitting a PPM [Hielscher et al., 2014]. They use a weighted $k$-nearest neighbour ($k$NN) classifier with Euclidean distance to measure similarity, and choose $k$ to be the value that results in the best average sensitivity, specificity, and accuracy after a two-fold validation. Another neighbourhood-based algorithm, proposed by Ng et al, uses a Locally Supervised Metric Learning (LSML) similarity metric, which is based on the Mahalanobis distance measure [Ng et al., 2015]. To assess the performance of their resulting PPM, they use the area under the receiver operating characteristic curve (AUROC), a commonly calculated measure of model discrimination. A more recent paper by Wang et al. also uses Mahalanobis distance to define similarity between patients in their neighbourhood-based algorithm [Wang et al., 2021]. They choose the optimal size of subpopulation based on the AUROC, cross-entropy loss and the F1-score, which is a measure of the model’s accuracy.
@@
Non-distance-based similarity metrics used in neighbourhood-based algorithms include a similarity measure proposed by Campillo et al. which uses the logical OR-operator [Campillo-Gimenez et al., 2013]. This measure incorporates weights that need to be learned before they are applied to the features. Then, the similarity measure is used to define a subpopulation consisting of the top $M$ similar patients to an index patient, where $M$ is chosen to maximize the AUROC. Another example is the cosine-similarity measure used in an algorithm proposed by Lee et al. (2015), where the measure is used to define a subpopulation consisting of the top $M$ similar patients to the index patient, where, again, $M$ is chosen to maximize the AUROC. In later work done by Lee, he suggests that choosing the optimal $M$ value will become important when taking these methods into practice [Lee, 2017].
@@
As mentioned above, a common quantity in choosing the subpopulation size $M$ is the AUROC [Sharafoddini et al., 2017, Allam et al., 2020], a measure of model discrimination. Discrimination is the ability of the predictive model to distinguish between outcome classes. Another measure of model performance is calibration, which is the agreement between the observed outcomes and the predictions. For prediction models in general, it has been criticized that calibration is not assessed nearly as often as discrimination [Van Calster et al., 2019]. This is an issue because poorly calibrated clinical algorithms can be misleading and potentially harmful when making decisions. Van Calster et al. also argue that even when calibration is assessed, it is usually done via external validation after the model has been developed via internal validation. Motivated by this issue, Jiang et al. propose a prediction algorithm that jointly optimizes the discrimination and calibration of a support vector machine by manipulating the decomposition of an overall performance measure, the Brier Score [Jiang et al., 2012]. The Brier Score is a summary measure that is optimized when the correct probabilities are used, making it a proper scoring rule [Gneiting and Raftery, 2007]. For models that have similar discrimination, a well-calibrated model has a better Brier Score than a miscalibrated model. However, in a recent paper by Assel et al., they claim that the Brier Score does not appropriately evaluate the clinical utility of prediction models [Assel et al., 2017].
@@
The remainder of the paper will be structured as follows. In Section 2, we propose an algorithm inspired by Lee et al. to develop a PPM that jointly optimizes both discrimination and calibration. We do this by finding the optimal size of the subpopulation, $M$, that jointly optimizes these performance measures through the use of a proposed loss function, which is an extension of the Brier Score. We use a two-term decomposition of the Brier Score, along with a mixture term, to allow for the flexibility of emphasizing one performance measure over another when training the PPM to find the optimal $M$. Section 3 contains results from a set of simulation studies which show the relationships between $M$ and various performance measures in addition to a sensitivity analysis on the mixture term of our defined loss function. In Section 4, our proposed algorithm is tested on data from the eICU Collaborative Research Database, a large multi-center critical care database from hospitals across the United States, to predict mortality of patients with diseases of the circulatory system [Pollard et al., 2018].
@@ 2.1 Notation
Let there be $n_{train}$ patients in the training data, where each patient $k$ is represented by a Euclidean vector, $\boldsymbol{X}_{k}$, $k$ = 1, …, $n_{train}$, in the multi-dimension predictor space defined by $p$ predictors. We let $x_{h}$, $h=1,...,p$, represent the $n_{train}\times 1$ vector consisting of each patient’s value of a given predictor. Further, let $\boldsymbol{X}_{I}$, $I=1,...,n_{test}$, be the Euclidean vector representation of the index patient, for whom we would like to make a prediction. Note that the index patient is not included in the training dataset.
The focus of the prediction task in this paper is on binary classification. Thus, we consider a binary outcome. The actual outcome for a patient $I$ in the testing dataset is represented by $y_{I}\in\{0,1\}$. Let $M$ be a scalar representing the size of the subpopulation in the training data consisting of similar patients to the index patient. Finally, let $p_{I}^{(M)}\in[0,1]$ represent the predicted probability of the outcome for a patient $I$ found using a model fit on a similar subpopulation of size $M$; we would add further specificity to $p_{I}^{(M)}$ in the case where there are competing models of the same and/or different type(s), such as $p_{I(logis1)}^{(M)}$, $p_{I(logis2)}^{(M)}$ or $p_{I(RF)}^{(M)}$ for two competing logistic regression models and a random forest model, respectively.
@@ 2.2 Patient Similarity Metric
We use the cosine similarity metric as it is bounded, interpretable, and able to handle non-continuous data. Further, it is hypothesized that the cosine metric may be more sensitive to details in patient differences when measuring similarities. [Sharafoddini et al., 2021] The cosine similarity between patients $I$ and $k$ is shown in Equation (1) where $\bullet$ and |||||$\cdot$||| | ⋅ | | represent the dot product and the Euclidean vector magnitude, respectively. We label this metric $CSM(\boldsymbol{X}_{I},\boldsymbol{X}_{k})$.
$CSM(\boldsymbol{X}_{I},\boldsymbol{X}_{k})=\frac{\boldsymbol{X}_{I}\bullet\boldsymbol{X}_{k}}{||\boldsymbol{X}_{I}||||\boldsymbol{X}_{k}||}$ (1)
Equation (1) is a bounded measure of similarity, with $-1\leq CSM(\boldsymbol{X}_{I},\boldsymbol{X}_{k})\leq 1$. $CSM(\boldsymbol{X}_{I},\boldsymbol{X}_{k})=-1$ corresponds to minimum (opposite) similarity, $CSM(\boldsymbol{X}_{I},\boldsymbol{X}_{k})=1$ corresponds to maximum similarity, and a value of 0 refers to orthogonal similarity.
@@ 2.3 Performance Measures
To measure the performance of a PPM, one should measure both the discrimination and calibration of the model [Steyerberg, 2019]. Often, overall performance measures such as the Brier Score, shown in Equation (2), are used, as the score is affected by both model discrimination and calibration.
$BrS=\frac{1}{n_{test}}\sum_{I=1}^{n_{test}}(y_{I}-p_{I}^{(M)})^{2}$ (2)
In a binary setting, such as the one we consider in this project, the Brier Score can range from 0 for a perfect model to 0.25 for a non-informative model when the outcome proportion is 50%. However, the range of the Brier Score changes depending on the outcome proportion, thus making the interpretation of the Brier Score a function of that proportion. A solution to this problem is to scale the Brier Score by its maximum possible score for a realized outcome proportion under a non-informative model [Steyerberg et al., 2010]. However, in this work, we do not consider the scaled Brier Score as the interpretation of the Brier Score is unnecessary for our purposes, as will be demonstrated below.
Even when predictive models have high model discrimination, the calibration may still be poor due to statistical overfitting [Van Calster et al., 2019]. One limitation of the Brier Score is that it may select a model that is miscalibrated with good discrimination over a model that is better calibrated with average discrimination [Assel et al., 2017]. Thus, it may be advantageous to evaluate the discrimination and calibration separately to obtain a more comprehensive assessment of the predictive performance of the model.
A common measure of discrimination is the area under the receiver operating characteristic curve (AUROC), which plots the sensitivity (true positive rate) against 1 – specificity (false positive rate) for consecutive cutoffs for the probability of an outcome. A close-to-perfect discriminated model will have an AUROC close to 1, while a model that cannot distinguish between classes will have an AUROC close to 0.5.
Van Calster et al. [Van Calster et al., 2016] have defined a hierarchy of four increasingly strict levels of calibration, referred to as mean, weak, moderate, and strong calibration. We will be using mean, weak, and moderate measures of calibration, as strong calibration is the utopic goal but unrealistic in practice.
Calibration-in-the-large (CITL) is a measure of mean calibration and evaluates whether the observed event rate is equal to the average predicted probabilities, shown in Equation (3). A CITL value of 0 indicates perfect mean calibration.
$CITL=\frac{1}{n_{test}}\sum_{I=1}^{n_{test}}|y_{I}-p_{I}^{(M)}|$ (3)
@@
The Integrated Calibration Index (ICI), valued between 0 and 1, is a measure of moderate calibration [Austin and Steyerberg, 2019], and is the weighted difference between observed and predicted probabilities, where the observations are weighted by the empirical density function of the predicted probabilities. This measure is used to compare the relative calibration of different prediction models, where a lower value of the ICI corresponds to better calibration. An estimate of the ICI is shown in Equation 4, where $\hat{p_{I}}^{(M)}$ are the smoothed predicted probabilities based on the LOESS calibration curve. This estimate of the ICI intrinsically includes the weights derived from the empirical density function of the predicted probabilities.
$ICI=\frac{1}{n_{test}}\sum_{I=1}^{n_{test}}\left|\hat{p_{I}}^{(M)}-p_{I}^{(M)}\right|$ (4)
@@ 2.4 Mixture Loss Function
$L^{(M)}=\frac{\alpha}{n_{test}}\sum_{I=1}^{n_{test}}(y_{I}-p_{I}^{(M)})(1-2p_{I}^{(M)})+\frac{1-\alpha}{n_{test}}\sum_{I=1}^{n_{test}}p_{I}^{(M)}(1-p_{I}^{(M)}).$ (5)
The first and second terms after the equal sign are related to calibration and discrimination, respectively. Specifically, the first term measures the lack of calibration, with an expectation of 0 under perfect calibration. The second term measures the lack of spread of the predictions. For larger values of $\alpha$, greater emphasis is put on calibration, while for smaller values of $\alpha$, greater emphasis is placed on discrimination. When $\alpha=0.5$, the loss function is proportional to the Brier Score.
@@ 2.5 Personalized Predictive Modelling Algorithm
We propose an algorithm which sets up a validation procedure to tune the subpopulation size, $M$, for a given value of $\alpha$. We include a hold-out validation step in the algorithm where we estimate the performance of the resulting PPM along with quantifying the uncertainty of these estimates. Step-by-step details of the algorithm are given below.
@@
      Randomly split the TrTe set into distinct training and testing sets using $K$-fold cross-validation (CV).
@@
        For index patient $I$ in the test set, calculate $d(\boldsymbol{X}_{I},\boldsymbol{X}_{k})$, for $k=1,...,n_{train}$ in the training set, where $d(\boldsymbol{X}_{I},\boldsymbol{X}_{k})$ is the similarity metric between the index patient and the $k^{th}$ training set patient.
@@
          Sort all $d(\boldsymbol{X}_{I},\boldsymbol{X}_{k})$ in descending order, and the patients corresponding to the top $M$ scores form the subpopulation.
@@
          Apply a pre-determined weighting function (eg. uniform, (half-)tri-cube) to each patient in the subpopulation. Note that the (half-)tri-cube weights give those more similar to the index patients a greater weight.
@@
      Draw a bootstrap sample from the validation sample (each the same size as the validation sample).
@@
        For index patient $I$ in the test set, calculate $d(\boldsymbol{X}_{I},\boldsymbol{X}_{k})$, for $k=1,...,n_{train}$ in the training set, where $d(\boldsymbol{X}_{I},\boldsymbol{X}_{k})$ is the similarity metric between the index patient and the $k^{th}$ patient.
@@
    Repeat the entire process $Z$ times, ie. repeat the process with $Z$ different validation holdout samples (and, thus, TrTe samples) to view the consistency of the tuning of $M$ (step (1) above) as well as the model performance measures in different hold-out validation samples. We suggest $Z$ to be approximately 10.
The proposed algorithm above is done for a given $\alpha$, perhaps one that is determined ahead of time, such as $\alpha=0.5$, where both discrimination and calibration are equally balanced in Equation (5). However, there may be some situations where we want to investigate the performance and resulting $M$ for different scenarios, say a bit more emphasis being placed on calibration over discrimination. We could then implement the above algorithm for a few different $\alpha$ values, perhaps to see how the performance will do as a function of both $\alpha$ and $M$. We will see some of this work in Sections 3 and 4 below.